OpenStudy (anonymous):

\[\int\limits_{-1}^{0}{1 \over 2x-1}dx\] Help with steps of integration and substitution.

OpenStudy (mikey):

Try a u substitution for 2x-1.

OpenStudy (anonymous):

I first need to integrate it.

OpenStudy (mikey):

That's what you do with u substitution it allows to write everything in terms of u and du so it makes it easier to integrate it.

OpenStudy (mikey):

u=2x-1...du = 2dx.

OpenStudy (anonymous):

Oh, yes... But I'm not sure how to do that with lns

OpenStudy (mikey):

Ok...so then you can rewrite the whole equation as 1/u (1/2du). Then pull the constant 1/2 out of the integral and integrate 1/u du.

OpenStudy (anonymous):

Not sure how to do that either..?

OpenStudy (mikey):

Does the first part of substituting u for 2x-1 make sense?

OpenStudy (anonymous):

Yes

OpenStudy (mikey):

And you know how we found what dx equals?

OpenStudy (diyadiya):

\[\int\limits \frac{1}{2x-1}dx\] u=2x-1 \[\frac{du}{dx}= \frac{d}{dx}(2x-1)= 2\] \[du=2dx\]

OpenStudy (anonymous):

Yes, I understand... but how do you integrate 1/u du.

OpenStudy (mikey):

You just integrate everything in terms of u instead of x. So the integral of 1/u is ln(u).

OpenStudy (anonymous):

And what does the "u" represent?

OpenStudy (diyadiya):

u=2x-1 we'll substitute 2x-1 later

OpenStudy (anonymous):

Oh, Ok.

OpenStudy (mikey):

2x-1. So you can either resubstitute 2x-1 back in for u at this point and use your original limits of integration. Or you can plug your original limits into the u=2x-1 equation and use those new values in the ln(u) equation.

OpenStudy (anonymous):

Can you show the steps of the substitution?

OpenStudy (anonymous):

I'm not sure I know what the final equation is before I substitute?

OpenStudy (mikey):

Alright, you start with 1/2x-1. Then you let u =2x-1. Then you need to find dx by taking the derivative of 2x-1. So when you plug the u in for 2x-1 and 1/2 du in for dx...that is your new integrand.....1/u times 1/2du.

OpenStudy (mikey):

And then you just integrate 1/2(1/u)du...which gives you the antiderivative of ln(u).

OpenStudy (anonymous):

How do you do that?

OpenStudy (mikey):

Do what exactly?

OpenStudy (anonymous):

integrate 1/2(1/u)du

OpenStudy (anonymous):

\[\int\limits_{-1}^{0} 1/(2x-1)\] let u=2x-1 du=2dx when x=-1, u=-3 when x=0, u=-1 now limit is from -3 to -1 = \[\int\limits_{-3}^{-1}(2/u) du\] =2 ln(-3)-ln(-1) =2 ln(-3/-1) because lna-lnb =ln(a/b) =2 ln(3) =\[\ln(3^{2})\] = ln(9) =2.1972

OpenStudy (mikey):

The 1/2 is a constant right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to write the answer here is really difficult

OpenStudy (mikey):

And you know that constants can pulled out of the integrand...right?

OpenStudy (anonymous):

Yes

OpenStudy (mikey):

Great. So we pull it out...and all that's left to integrate is 1/u du. Do you know how to integrate 1/x dx?

OpenStudy (anonymous):

It's Inx But doesn't u stand for 2x-1

OpenStudy (anonymous):

integrate (1/x dx)

OpenStudy (mikey):

Forget about that for a second...because we're not integrating for x anymore...we're integrating for u.

OpenStudy (anonymous):

So, it would be lnU

OpenStudy (mikey):

Right! Now....what does U equal?

OpenStudy (anonymous):

2x-1

OpenStudy (mikey):

So you can now put that back in for u in the lnu equation....but don't forget about the 1/2 we took out...which gets multiplied with the ln function.

OpenStudy (anonymous):

Ok, so it's 1/2 * In(-1) - In(-3)

OpenStudy (mikey):

Yes...1/2ln(-1)-1/2ln(-3)

OpenStudy (anonymous):

Ok, so the answer is -1/2ln3

OpenStudy (mikey):

Except you cant take logs of negative numbers!

OpenStudy (anonymous):

Yes, that's the answer in the book :)

OpenStudy (anonymous):

Thanks for helping me!

OpenStudy (mikey):

Ok....if they're cool with that. You bet. :)

OpenStudy (anonymous):

Yes, they are :)